A body goes from a point A to another point B 150 m away in 30 second and return back in 20 sec. Then the average velocity is

Grade:9

1 Answers

utpal raj
7 years ago
@ vidith 
 on going from A to B distance travelled =150m
t1= 30 sec
so v1=150/30=5m/s
v2=150/20=7.5m/s
avg velocity=(v1+v2)/2
so avg velocity=6.25m/s
hope you understand 
good luck....